OpenStudy (anonymous):

What is 4x + 2y = -20 written in slope-intercept form

OpenStudy (anonymous):

y = -2x - 10

OpenStudy (anonymous):

y = -2(x + 5)

OpenStudy (anonymous):

thanks rajapurna for the help

OpenStudy (anonymous):

nooo x /a + y/b = 1 is intercept form

OpenStudy (anonymous):

What is the slope of the graph for the equation y = -5x + 8

OpenStudy (anonymous):

so, \[\frac{ x }{ -5} + \frac{ y }{-10 } = 1\]

OpenStudy (anonymous):

- 5 is the Slope of y = -5x + 8

OpenStudy (anonymous):

thanks Write the slope-intercept form of the equation of the line that passes through (4, 2) with a slope of 3

OpenStudy (anonymous):

answers are y = 3x - 10 y = 3x + 4 y - 2 = 3 (x - 4) y = 4x + 3